### v3

# DriveIndia: An Object Detection Dataset for Diverse Indian Traffic Scenes

Abstract
We introduce DriveIndia, a large-scale object detection dataset purpose-built to capture the complexity and unpredictability of Indian traffic environments. The dataset contains 66,986 high-resolution images annotated in YOLO format across 24 traffic-relevant object categories, encompassing diverse conditions such as varied weather (fog, rain), illumination changes, heterogeneous road infrastructure, and dense, mixed traffic patterns and collected over 120+ hours and covering 3,400+ kilometers across urban, rural, and highway routes. DriveIndia offers a comprehensive benchmark for real-world autonomous driving challenges. We provide baseline results using state-of-the-art YOLO family models, with the top-performing variant achieving $mAP_{50}$ of 78.7%. Designed to support research in robust, generalizable object detection under uncertain road conditions, DriveIndia will be publicly available via the TiHAN- IIT Hyderabad dataset repository (https://tihan.iith.ac.in/TiAND.html).
I Introduction
Autonomous driving systems rely heavily on robust object detection to perceive and navigate complex environments. While significant progress has been made using deep learning, particularly convolutional neural networks (CNNs), most public benchmarks such as KITTI [1] and nuScenes [2] are designed around structured, rule-compliant traffic scenarios common in Western countries. However, these conditions differ dramatically from those encountered in regions like India, where unstructured traffic, variable infrastructure, and environmental uncertainty are the norm.
In India, road conditions are uniquely challenging. Urban and rural areas often lack standardized signage or lane markings, while traffic comprises a heterogeneous mix of vehicles, pedestrians, animals, and non-motorized transport. These factors make India an ideal testbed for developing generalizable perception models that can handle uncertainty and real-world complexity.
Several Indian datasets have emerged to address local challenges, such as the Indian Driving Dataset (IDD) [3, 4], DATS [5], NITCAD [6], and the recent multimodal TIAND dataset [7]. While these provide valuable resources, they often focus on specific tasks (e.g., semantic segmentation), limited object classes, or urban-centered use cases, leaving a gap in large-scale, diverse object detection datasets covering varied road types and environmental conditions.
To address this gap, we introduce DriveIndia (refer Fig. 1), a large-scale object detection dataset explicitly tailored for Indian road environments. It contains 66,986 high-resolution RGB images annotated in the YOLO format, with normalized bounding boxes stored in per-image .txt files. The annotations span 24 traffic-relevant object classes. The data was captured under diverse conditions—including fog, rain, low light, dense traffic, and infrastructure anomalies—making DriveIndia well-suited for training models that need to perform reliably in unstructured and unpredictable real-world scenarios. All images are recorded at a resolution of 1920×1080 pixels, ensuring clear visual quality for object localization and model training.
While prior work such as TIAND [7] focuses on multimodal autonomy using LiDAR and radar sensors, DriveIndia emphasizes general-purpose visual object detection across geographic, environmental, and traffic diversity. It further provides a benchmark for evaluating real-time detectors using YOLO-series models [8], establishing new baselines for Indian road conditions.
Figure 1: Sample scenes from the DriveIndia dataset: urban roads, rural highways, fog, and rain.
Key Contributions
  • Developed DriveIndia, the largest publicly available object detection dataset focused on unstructured Indian traffic, featuring $\sim$67,000 annotated images.
  • It covers 24 traffic-relevant object classes labeled in YOLO format with normalized bounding boxes.
  • We establish strong detection baselines using state-of-the-art models.
  • The dataset will be publicly released via the TiHAN-IIT Hyderabad portal [9] to support open research.
TABLE I: Comparison of major Indian and global object detection datasets. DriveIndia offers the largest volume of annotated images under uncertain weather conditions (fog, rain), while IDD and DATS focus on structured urban scenes.
Attribute DriveIndia (Ours) IDD DATS ITD BDD100K
Release Year 2025 2023 2022 2020 2020
Task Type Object Detection 3D Object Detection Object Detection Vehicle Classification Detection / Segm. / Tracking
Images 66,986 46,588 10,000+ 17,666 (9.8k from IDD) 100,000
Object Classes 24 17 (10 used) 45  10 10
Camera Type Vehicle-mounted Vehicle-mounted Vehicle-mounted Fixed (CCTV) Vehicle-mounted
Environment Urban, Rural, Highway Urban, Suburban Urban, Rural Junction-level All (urban, rural)
Weather Conditions Fog, Rain Clear only Clear only Clear only Fog, Rain, Night
Night-time Data No No No No Yes
Annotation Format TXT (YOLO-style) COCO (JSON) XML / TXT / JSON Class Labels COCO / MOT / Video
$\mathbf{mAP_{50}}$ 78.7% 67.5% Not Reported 53.9% 56.7%
$\mathbf{mAP_{50{:}95}}$ 51.7% 33% 31% 42% 39%
We hope DriveIndia is a foundational resource for advancing perception models in diverse, real-world traffic environments, particularly for emerging markets.
II Related Work and Dataset Comparison
Object detection in real-world driving scenarios has advanced rapidly with the release of large-scale datasets. However, most public benchmarks such as KITTI, Cityscapes [10], and BDD100K [11] are designed around structured, rule-compliant traffic environments typically found in developed countries. These datasets, while valuable, fall short of representing the unstructured, heterogeneous, and often chaotic nature of traffic conditions prevalent in countries like India.
In recent years, several Indian datasets have emerged to address this gap. The Indian Driving Dataset (IDD) [3], initially released in 2018, focuses on semantic segmentation in unstructured roads, particularly in urban and suburban environments. A more recent extension, IDD [4], includes bounding box annotations for object detection across 46,588 images, but does not cover adverse weather conditions.
DATS [5] was curated to support object detection in dense urban Indian traffic and includes critical classes such as autorickshaws and pedestrians. Although it annotates in XML, TXT, and COCO formats, it is geographically constrained and lacks diversity in environmental settings.
The Indian Traffic Dataset (ITD) [12] from IIT Roorkee combines 9,888 images from IDD with 7,778 newly collected ones, focusing on vehicle classification for adaptive traffic control. It uses fixed surveillance-style camera setups and does not provide detailed bounding box annotations suitable for mobile perception systems.
For global comparison, BDD100K [11] remains the most comprehensive benchmark with 1,00,000 images covering object detection, semantic segmentation, and tracking across varied weather, time-of-day, and road conditions. While it includes night-time and rainy scenarios, it lacks the localized diversity of road users, signage, and informal traffic behavior in Indian settings.
In addition to dataset-focused efforts, several studies have explored the performance of deep learning models on Indian road scenes. Mukhopadhyay et al. [13] presented a comparative analysis of different CNN architectures for object detection on localized road datasets, highlighting model selection and feature extraction challenges. Similarly, Sehgal et al. [14] conducted a comparative study of deep learning models for vehicle detection in unconstrained Indian road environments, underscoring the performance limitations of standard architectures in chaotic, heterogeneous traffic conditions. While these works provide valuable insights into model behavior, they primarily focus on architectural comparisons rather than addressing the broader need for large-scale, diverse datasets covering multiple object classes and environmental conditions, which DriveIndia aims to fulfill.
Compared to these, DriveIndia offers a significantly larger and more diverse set of object annotations, covering challenging scenarios such as fog, rain, and unstructured rural roads. It enables benchmarking of real-time object detectors under conditions typical of developing regions. A detailed comparison is provided in Table I.
III DriveIndia Dataset Overview
DriveIndia is a large-scale object detection dataset created to reflect the complexity and uncertainty of real-world traffic conditions in India. It consists of 66,986 high-resolution RGB images, split into 53,586 for training, 6,700 for validation, and 6,700 for testing, captured over 120+ hours of driving and covering more than 3,400 kilometers across highways, urban centers, and rural regions. The dataset comprises a total of approximately 4,71,092 annotated instances across 24 traffic-relevant object categories. Data was recorded using vehicle-mounted cameras operating at 1920×1080 resolution, providing detailed visual context for object detection tasks.
This section outlines the data collection methodology, annotation process, and environmental diversity represented in the dataset.
III-A Data Collection Strategy
We collected data across a wide geographic spread in southern India, spanning highways, city streets, rural roads, and construction areas beyond Hyderabad. The recording setup used multiple synchronized RGB cameras mounted on passenger vehicles (see Fig. 2).
Figure 2: Data Collection Vehicle with Front & Back Camera
The collection periods were carefully chosen to span different times of day and to include a wide range of environmental scenarios.
III-B Annotation Format and Tools
All images in DriveIndia were annotated using the YOLO format, where each object instance is represented as a single line in a per-image ‘.txt’ file.
Annotation was performed using a combination of LabelImg [15] and custom validation scripts to ensure consistency and accuracy.
The 24 object classes included in DriveIndia are grouped as follows see Fig. 3 for their frequency distribution:
  • Person : Pedestrian
  • Vehicles: Bicycle, Car, Motorcycle, Bus, Commercial vehicle, Truck, Autorickshaw, Ambulance, Police vehicle, Tractor, Pushcart, Construction vehicle
  • Road Infrastructure: Route board, Traffic sign, Traffic light, Temporary traffic barrier, Traffic cone, Rumblestrips, Unmarked speed bump, Marked speed bump, Zebra crossing
  • Environmental & Surface Anomalies: Animal, Pothole
These categories were selected to reflect the vast diversity of road users and objects commonly encountered in Indian traffic environments, including region-specific classes not present in most international datasets.
Figure 3: Distribution of class labels in the DriveIndia Dataset
III-C Weather and Scene Diversity
One of DriveIndia’s distinguishing features is its deliberate inclusion of challenging environmental conditions rarely addressed in mainstream datasets. These include:
  • Weather: Fog, rain, overcast skies, and varying visibility levels.
  • Lighting: Low-light and uneven lighting conditions (e.g., shadowed roads).
  • Scenes: Urban congestion, rural roads, flyovers, intersections, roadwork zones, and unmarked lanes.
This diversity is essential for developing object detection models that generalize across unpredictable and complex traffic environments, especially in developing regions.
III-D Data Annotation Guidelines
Annotations in the DriveIndia dataset were performed using strict, evidence-based protocols to ensure high-quality and consistent labeling. The key guidelines followed are:
  • Bounding Box Policy:
    • Boxes were drawn tightly around the main object, excluding small extremities and accessories.
    • Occluded objects were annotated only for the visible parts, with emphasis on key regions (e.g., heads/shoulders for humans, wheels for cycles).
    • Riders were annotated separately from the vehicles they operate.
    • Edge-cut or motion-blurred objects were labeled only if at least 50% of the object was visible.
  • Image Filtering:
    • Images with severe blur were filtered from validation and test sets.
    • A Laplacian variance threshold of 20 was used to exclude low-clarity images [16].
  • Quality Assurance:
    • Two-stage verification: random sampling and senior review for annotation consistency.
    • Automated checks ensured:
        Each image had a valid corresponding label file.
        Empty or invalid annotations were removed.
        Class distributions were monitored across train/test/val splits to avoid data leakage.
  • Design Motivation:
    • The guidelines were informed by recent research showing that tight and clean annotations directly improve object detection performance [17].
Figure 4: Example detections from the DriveIndia dataset, showcasing diverse object categories including vehicles, pedestrians, infrastructure elements, and environmental obstacles under varied conditions.
III-E Dataset Organization
The DriveIndia dataset is organized into standard training, validation, and test splits. The directory structure, shown below, separates images from their corresponding YOLO format annotation files.
[…]
.1 DriveIndia/. .2 images/. .3 train/. .3 val/. .3 test/. .2 labels/. .3 train/. .3 val/. .3 test/.
This organization ensures compatibility with standard object detection frameworks. Each image file has a corresponding .txt annotation file in the matching subdirectory.
IV Benchmark: Object Detection
To establish DriveIndia as a robust benchmark for object detection in uncertain real-world driving conditions, we evaluated several recent, high-performing detection models—including multiple YOLO versions [8, 18, 19, 20, 21] and state-of-the-art transformer-based detectors such as RT-DETR [22] and EfficientDet [23]. This section presents baseline results and discusses challenges uncovered by class-wise evaluation.
IV-A Experimental Protocol
All models were trained on the DriveIndia dataset with a standardized protocol: 54,856 images for training and 6,700 each for validation and testing. We adopted common hyperparameters across all runs for fairness. Model performance was assessed on the test set using widely adopted detection metrics: $mAP_{50}$, $mAP_{50{:}95}$, precision, and recall.
IV-B Quantitative Results
Table LABEL:tab:yolo_results summarizes the results for all evaluated models. The YOLO family remains highly competitive, with YOLOv8 achieving the highest $mAP_{50}$ (78.7%), while YOLOv9 offers the best precision/recall balance. Figure  4 visualizes typical DriveIndia scenes by the YOLOv8 Model, reflecting the dataset’s real-world diversity.
IV-C Per-Class Performance Analysis
A detailed breakdown of class-wise metrics is presented in Table II. While common objects like car, truck, and bus achieve high $mAP_{50}$ scores (>0.7>0.7), rare and small-scale classes such as pushcart, pothole, police vehicle, and unmarked speed bump show lower detection accuracy ($mAP_{50{:}95}<0.5$). Figure 5 displays the confusion matrix, highlighting frequent misclassifications, particularly among visually similar or underrepresented classes.
TABLE II: Per-class detection performance on the DriveIndia validation set. Color indicates high, moderate, and low detection scores.
Class Images Instances $\mathbf{mAP_{50}}$ $\mathbf{mAP_{50{:}95}}$ Pedestrian 4,344 11,806 0.754 0.501 Bicycle 73 80 0.645 0.455 Car 4,911 11,959 0.945 0.819 Motorcycle 4,156 9,795 0.947 0.703 Route board 298 333 0.841 0.587 Bus 997 1,181 0.878 0.735 Commercial vehicle 1,457 1,798 0.840 0.689 Truck 2,036 2,918 0.899 0.752 Traffic sign 684 873 0.851 0.565 Traffic light 297 432 0.710 0.353 Auto-rickshaw 2,291 3,242 0.940 0.803 Ambulance 20 20 0.780 0.659 Construction vehicle 149 156 0.735 0.515 Animal 62 135 0.769 0.471 Unmarked speed bump 37 41 0.454 0.182 Marked speed bump 78 91 0.715 0.343 Pothole 189 358 0.497 0.232 Police vehicle 10 10 0.511 0.470 Tractor 173 178 0.843 0.633 Pushcart 26 29 0.391 0.206 Temporary traffic barrier 145 278 0.837 0.547 Rumble strips 355 431 0.825 0.488 Traffic cone 181 567 0.800 0.367 Zebra crossing 159 196 0.684 0.339
Figure 5: Confusion Matrix, highlighting confusion between visually or contextually similar categories.
IV-D Discussion: Class-Specific Detection Challenges
Our evaluation [^1] Color categories for Table II For $mAP_{50:95}$, green >>0.50, orange 0.30–0.50, red <<0.30. For $mAP_{50}$, green $\geq$0.75, orange 0.50–0.75, red <<0.50. as seen on Table II shows that common road users—such as car, Auto-rickshaw, truck, and motorcycle—are detected with high accuracy ($mAP_{50}$ above 0.9), reflecting model robustness for prevalent classes in Indian traffic.
However, detection is notably weaker for rare or visually ambiguous classes. Bicycle exhibits lower performance ($mAP_{50}$ 0.645), likely due to frequent occlusion and visual overlap with motorcycles. Similarly, classes like animal are often confused or missed due to appearance variability and limited samples.
Critical but underrepresented categories—such as ambulance ($mAP_{50}$ 0.78) and police vehicle ($mAP_{50}$ 0.51)—also pose challenges, as these often lack distinct visual features and are seen infrequently. Infrastructure-related classes such as pothole and unmarked speed bump achieve the lowest mAP, underscoring the difficulty of detecting small or subtle road features.
These results emphasize the need for improved data augmentation and targeted model enhancements for ambiguous and rare, yet safety-critical, classes—key priorities for robust ADAS in real-world Indian traffic scenarios.
V Dataset Availability and Licensing
The DriveIndia dataset will be publicly released for academic and non-commercial research via the official TiHAN-IIT Hyderabad dataset portal:
https://tihan.iith.ac.in/tiand-datasets/
The release will include:
  • 66,986 RGB images annotated in YOLO format
  • 24 traffic-relevant object categories
  • Pre-defined training, validation, and test splits
VI Conclusion
We presented DriveIndia, a comprehensive object detection benchmark designed to reflect the full complexity and diversity of Indian traffic environments. With nearly 67,000 annotated images spanning 24 object categories—including rare and safety-critical classes—DriveIndia uniquely captures the real-world challenges of heterogeneous road users, occlusions, and adverse weather conditions such as fog and rain.
DriveIndia is fully compatible with modern detection frameworks, including YOLOv5, YOLOv8, YOLOv9, YOLOv11, RT-DETR, and EfficientDET, facilitating rigorous comparisons across architectures. Our large-scale experiments reveal that even state-of-the-art detectors face significant drops in accuracy for rare, small, or ambiguous classes (e.g., pushcart, unmarked speed bump, bicycle), despite strong overall $mAP$ on common vehicles.
These results highlight persistent gaps in current object detection approaches and the importance of context-aware models, targeted data augmentation, and improved handling of long-tail categories. By releasing DriveIndia, we aim to accelerate research towards safer, more reliable ADAS and autonomous systems in unstructured, real-world scenarios.
We invite the research community to leverage and extend DriveIndia, and we welcome contributions that push the boundaries of robust visual perception in challenging traffic environments.
VII Limitations and Future Research
Despite the scale and diversity of DriveIndia, certain limitations remain that present opportunities for further advancement:
  • Geographical Scope: The dataset primarily covers southern Indian regions and does not capture the full diversity of terrains, traffic patterns, and signage found across other parts of India, such as hilly terrains or dense metro areas in the north and northeast.
  • Sparse Representation of Rare Classes: Safety-critical but rare classes (e.g., ambulance, police vehicle have limited instances, challenging robust model training for long-tail categories.
  • Absence of Night-Time Scenes: Although varied weather and lighting conditions are included, night-time scenarios are underrepresented, affecting generalization in low-light settings.
  • Single-Modality Limitation: DriveIndia uses only RGB imagery. Incorporating additional modalities such as LiDAR would improve perception under adverse weather and complex road geometries.
  • Lack of 3D and Semantic Annotations: The dataset includes only 2D bounding boxes. Adding semantic masks and 3D bounding boxes (or point-level LiDAR) would enable broader AV perception tasks.
  • Benchmark Scope: Current evaluation is limited to object detection. Extending to tracking, segmentation, and domain adaptation would enhance research applicability for ADAS systems.
We envision DriveIndia as a foundational benchmark for robust, multimodal perception research on unstructured roads. Community contributions in these directions are encouraged.
Acknowledgements
This research was supported by the DST National Mission on Interdisciplinary Cyber-Physical Systems (NM-ICPS) and TiHAN at IIT Hyderabad, whose funding and resources were instrumental.
I express my sincere gratitude to Aravind Kumapati and Menasath Santhosh for their diligent efforts in data collection and annotation, which ensured the development of high-quality datasets critical to this work. I also extend my thanks to the anonymous team members whose contributions to data processing and annotation were invaluable.